OpenStudy (anonymous):

Please help :( solve the following equation with rational exponents and radicals (x-2)^2/3=49

OpenStudy (anonymous):

this website helps u a lot

OpenStudy (anonymous):

http://www.tiger-algebra.com/drill/%28x-2%29~2/3=49/

OpenStudy (darkigloo):

is that \[\frac{ (x-2)^2 }{ 3 } = 49\] ?

OpenStudy (anonymous):

no its( x-2) raised to the 2 thirds power equals 49

OpenStudy (darkigloo):

\[(x-2)^{2/3}=49\] multiply both sides by the inverse of 2/3 to cancel out the exponent on the left side. \[((x-2)^{2/3})^{3/2} = 49^{3/2}\] Then you'll have x-2 = 49^(3/2)

OpenStudy (anonymous):

k thanks